OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17 Someone please help, I have no clue what I am doing.

OpenStudy (anonymous):

A perpendicular line has a negative reciprocal for a slope.

OpenStudy (anonymous):

OK, I learned this, lol I just don't understand how to get the answer. The way everyone is explaining it, I just don't get it.

OpenStudy (bahrom7893):

3x+y=7, can you solve for y?

OpenStudy (bahrom7893):

Come on, subtract 3x from both sides, what do you get?

OpenStudy (anonymous):

Lol I am just viewing this, my baby cousin needed help. Ok let me work this out.

OpenStudy (anonymous):

X + Y = 4?

OpenStudy (anonymous):

Ok, I am positive I messed that whole thing up.

OpenStudy (anonymous):

Are you familiar with the point-slope formula? http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

Yes, I am, I am learning it in virtual school now. I just don't get it.

OpenStudy (anonymous):

I will take a look at that link

OpenStudy (anonymous):

You pick a point and substitute those values for \(x_1\) and \(y_1\). You should still have both and x and a y in your equation.

OpenStudy (anonymous):

Which numbers? I don't understand.

OpenStudy (anonymous):

Line: 3x + y = 7 Point: (6, −1) Point-Slope formula: \(y – y_1 = m(x – x_1)\)

OpenStudy (anonymous):

Because there is already an equation up there, and the coordinates, at this point I have no idea what to do with anything. I am so sorry.

OpenStudy (anonymous):

First, find the slope of the perpendicular line. What is the slope of the given line?

OpenStudy (anonymous):

To find the slope, solve for y. y=mx+b

OpenStudy (anonymous):

I do not know the numbers I am substituting, sorry if I was unclear.

OpenStudy (anonymous):

OK. The point (6,-1) are the numbers you will be substituting for \((x_1,y_1)\).

OpenStudy (anonymous):

\(x_1=6, y_1=-1\)

OpenStudy (anonymous):

what would 'm' be?

OpenStudy (anonymous):

m is the slope so it will be the negative reciprocal of the original slope.

OpenStudy (anonymous):

I have no clue what that number is. At all, I could most likely figure out the equation if I knew it. But I am having a hard time.

OpenStudy (anonymous):

Believe it or not, I am actually trying lol, I know I am bad at this.

OpenStudy (anonymous):

Start with the given equation: \(3x + y = 7\) and solve for y.

OpenStudy (anonymous):

I know you are trying. I'm here to help.

OpenStudy (anonymous):

Ok, would I take three to each sides and subtract?

OpenStudy (anonymous):

Subtract 3x from each side, yes.

OpenStudy (anonymous):

And you are not bad at this, you just haven't learned it yet. Give yourself time.

OpenStudy (anonymous):

So would that be Y + -3 =7 or would it be x + y = 4?

OpenStudy (anonymous):

\(3x + y = 7\) \(-3x~~~=-3x\) -------------- \(y=-3x+7\) Does that make sense?

OpenStudy (anonymous):

Yes, that makes much more sense. Ok, now I think I should divide now, but I don't know by what.

OpenStudy (anonymous):

If y had a coefficient, you would divide. Since it doesn't, you don't need to divide. Do you know what the slope of this line is?

OpenStudy (anonymous):

No, I can't say that I do.

OpenStudy (anonymous):

y=mx+b m=slope Does that help?

OpenStudy (anonymous):

I know m= slope, but I don't know what I am suppose to do now, I don't know where to carry the problem off to.

OpenStudy (anonymous):

Going back to the point-slope formula, substitute the negative reciprocal for m in that formula.

OpenStudy (anonymous):

I have no clue what that means lol.

OpenStudy (anonymous):

The reciprocal means to flip the fraction. Remember that \(3=\frac{3}{1}\)

OpenStudy (anonymous):

So three over 1? Can you show me what I have to do next?

OpenStudy (anonymous):

The reciprocal of \(3 = \frac{1}{3}\). The negative reciprocal of \(3 = -\frac{1}{3}\)

OpenStudy (anonymous):

I am still stuck.

OpenStudy (anonymous):

In summary: "Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1)." 1. Solve the given line for y. 2. Identify the slope of the given line. 3. Use the point-slope formula to determine the equation of the perpendicular line. 4. Substitute the x and y values from the given point into the point-slope formula. 5. Substitute the negative reciprocal for m in the point-slope formula. 6. Solve for y if requested.

OpenStudy (anonymous):

Does this help at all? Which step do you find confusing?

OpenStudy (anonymous):

Four, five, and six.

OpenStudy (anonymous):

I have to have this test turned in by eleven. But I do not want to make a bad grade on it :/ even though I am bad at math, I manage to make good grades, due to the fact that try extra hard for them.

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line \(3x + y = 7\) and passes through the point (6, −1). 1. \(y=-3x+7\) 2. \(m=-3\) 3. \(y – y_1 = m(x – x_1)\) 4. Using the point (6,-1). \(y – -1 = m(x – 6)\) Is this what you have?

OpenStudy (anonymous):

I don't believe anyone is "bad at math". If you put in the effort (which you are clearly doing) and find the best explanations for you, you will learn what you need to learn.

OpenStudy (anonymous):

Don't be afraid to use many different sources to figure things out. And feel free to play around. Math is all about pattern recognition.

OpenStudy (anonymous):

Did you have anything close to what I wrote?

OpenStudy (anonymous):

yes I got that, but I don't think that is a choice D:

OpenStudy (anonymous):

There are still more steps. In the original equation m=-3, in this equation, the slope should be the negative reciprocal. \(-3 \implies \frac{1}{3}\). Plug that value in for m in this equation.

OpenStudy (anonymous):

So, y-1= -3 ( x-6)

OpenStudy (anonymous):

And I need to solve?

OpenStudy (anonymous):

Your slope is the original slope. The slope for the perpendicular line needs to be the negative reciprocal.

OpenStudy (anonymous):

What is the negative reciprocal?

OpenStudy (anonymous):

negative means to negate the number. The negation of 4 is -4; the negation of -5 is 5. Negating a number is the same as multiplying it by -1.

OpenStudy (anonymous):

The reciprocal means to flip the fraction. Remember that a whole number can be represented by the whole number over 1. For example \(4 = \frac{4}{1}\).

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

So "m" would be Negative 4/1

OpenStudy (anonymous):

Your original slope is -3. So, multiply by -1 then flip the fraction.

OpenStudy (anonymous):

3/1?

OpenStudy (anonymous):

Now, flip the fraction.

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

So 1/3 is 'M"?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\(y–−1=\frac{1}{3}(x–6)\) Now solve the equation.

OpenStudy (anonymous):

Solve for y is what I meant.

OpenStudy (anonymous):

Y- (-1) right?

OpenStudy (anonymous):

If you mean \(y--1=y+1\) then yes. That is the first step to solving this equation.

OpenStudy (anonymous):

\(y+1=\frac{1}{3}(x–6)\)

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

I am asking if the one is a negative, I have no clue. I want to say you move some numbers around.

OpenStudy (anonymous):

Yes, y--1 = y-(-1).

OpenStudy (anonymous):

Can you explain to me how to solve this?

OpenStudy (anonymous):

Sure. You are familiar with the double negative making a positive? That is how y-(-1) = y+1.

OpenStudy (anonymous):

The next step would be to distribute the \(\frac{1}{3}\).

OpenStudy (anonymous):

What would that look like?

OpenStudy (anonymous):

Number wise.

OpenStudy (anonymous):

\(y+1=\frac{1}{3}(x–6)\) To distribute the \(\frac{1}{3}\) you need to multiply it by each of the terms inside the parenthesis. \(\frac{1}{3} \times x\) and \(\frac{1}{3}\times -6\)

OpenStudy (anonymous):

OK, let me try this, I will show you what I come out with.

OpenStudy (anonymous):

I would like to invite you to a whiteboard website. It might be easier to explain this.

OpenStudy (anonymous):

I can't do those things due to my limited gigabytes

OpenStudy (anonymous):

We can continue to work here. No problem.

OpenStudy (anonymous):

Ok So I came out with; y +1 = 1/3x -2

OpenStudy (anonymous):

Yes. One more step to finish.

OpenStudy (anonymous):

Ok, I assume it is to do something to the 1?

OpenStudy (anonymous):

You assume correctly.

OpenStudy (anonymous):

If I disappear for a bit it is because I am babysitting, and my little cousin is sick, so I am going to be up and down. So would we subtract.. orr?

OpenStudy (anonymous):

No problem. Yes, subtract 1 from both sides.

OpenStudy (anonymous):

So subtract one from itself and two?

OpenStudy (anonymous):

and the -2, yes.

OpenStudy (anonymous):

Ok let me try this.

OpenStudy (anonymous):

Effectively moving the 1 to the other side of the equation and combining it with the -2.

OpenStudy (anonymous):

Y= 1/3 x - 3 or Y= 1/3x + 3?

OpenStudy (anonymous):

\(y=\frac{1}{3}x-3\)

OpenStudy (anonymous):

So Y= 1/3x -3? :)

OpenStudy (anonymous):

yes, that is the correct answer.

OpenStudy (anonymous):

The answer is A? :D